OpenStudy (anonymous):

2 divided by ^6 to the 27. im new to tbis website so im not sure how to put it in properly.

OpenStudy (eg145341):

You could take a picture of the problem and attach it as a file. Or you could press the equation button, write it then post.

OpenStudy (anonymous):

\[2\div \sqrt[6]{27}\]

OpenStudy (marigirl):

@Ayahchehab \[ 2 \div \sqrt[6]{27}\] Let's begin with the term at the bottom which is \[\sqrt[6]{27}\] this is the same as saying \[27^{\frac{ 1 }{ 6 }}\] (you can check this using your calculator too, but you will notice that you get a decimal number as your answer, so we need to work around it) Next step now, There is no set rule, its more like simplifying and in my opinion, it comes with practice and experience/familiarity with maths. 27 is the same as 3^3 so now we can rewrite the bottom part as \[27^{\frac{ 1 }{ 6 }} = (3^{3})^{\frac{ 1 }{ 6}}\] Then we can multiply the powers across \[(3^{2})^{\frac{ 1 }{ 6 }}\] \[3^{\frac{ 3 }{ 6 }}\] simplify to \[3^{\frac{ 1 }{ 2 }}\] when there is a power of half, (like above) it is the same as the square root so it is the same as writing \[\sqrt{3}\] therefore the final answer is \[2\div \sqrt{3}\]